# Improving SDSS Cosmological Constraints through $\beta$-Skeleton Weighted Correlation Functions

### v1

@@ Abstract
The $\beta$-skeleton approach can be conveniently utilized to construct the cosmic web based on the spatial geometry distribution of galaxies, particularly in sparse samples. This method plays a key role in establishing the three-dimensional structure of the Universe and serves as a tool for quantitatively characterizing the nature of the cosmic web. This study is the first application of $\beta$-skeleton information as weights in mark weighted correlation functions (MCFs), presenting a novel statistical measure. We have applied the $\beta$-skeleton approach to the CMASS NGC galaxy samples from SDSS BOSS DR12 in the redshift interval $0.45\leq z\leq 0.55$. Additionally, we applied this approach to three COLA cosmological simulations with different settings ($\Omega_{m}=0.25,\Omega_{m}=0.31,\Omega_{m}=0.4$) for comparison. We measured three MCFs, each weighted by: i) the number of neighboring galaxies around each galaxy; ii) the average distance of each galaxy from its surrounding neighbors; iii) the reciprocal of the average distance of each galaxy from its surrounding neighbors. By comparing measurements and calculating corresponding $\chi^{2}$ statistics, we observe high sensitivity to the cosmological parameter $\Omega_{m}$ through a joint analysis of the two-point correlation and three MCFs.
@@ I Introduction
According to the principles of cosmology, the Universe is considered homogeneous and isotropic at a large scale. However, the combined influence of gravity and the ongoing expansion of the universe results in a complex and nonlinear structure. In terms of spatial configuration, the unique and intricate web structure formed by the anisotropic collapse of gravity is referred to as the “cosmic web” (Bardeen et al., 1986; de Lapparent et al., 1985; Huchra et al., 2012; Tegmark et al., 2004; Guzzo et al., 2014; Suárez-Pérez et al., 2021). Topologically, based on the dimensions of gravitational collapse, the cosmic web is typically categorized into knots, filaments, sheets, and voids.
The study of the structure of the cosmic web relies on large-scale redshift surveys, and currently, there are a number of large-scale observational projects, such as the Sloan Digital Sky Survey (SDSS) (York et al., 2000; Eisenstein et al., 2005; Percival et al., 2007; Anderson et al., 2012; Sánchez and ohers, 2012; Sánchez et al., 2013; Anderson et al., 2014; Ross et al., 2015; Beutler et al., 2017; Sánchez et al., 2017; Alam et al., 2017; Chuang et al., 2017), the 2-degree Field Galaxy Redshift Survey (2dFGRS) (Colless et al., 2001), the 6-degree Field Galaxy Redshift Survey (6dFGRS) (Beutler et al., 2012, 2011), and the WiggleZ Dark Energy Survey (Parkinson D., 2012).
The ongoing Stage-IV surveys, such as the Dark Energy Spectroscopic Instrument (DESI) DESI Collaboration et al. (2016, 2022), are expected to further deepen our understanding of the Universe. They possess the capability to observe wider and deeper fields of view with higher resolution, allowing for the collection of more accurate and extensive data and providing additional information about the non-linear scale structure of the universe.
@@
The exploration of the cosmic web has consistently been a significant topic in the study of the Universe. Since early maps of the Universe confirmed the existence of such webs through galaxy redshift measurements, there has been a continuous effort over the past 40 years to find a consistent and stable method to define these web-like structures. Currently, the commonly used analytical tools for studying cosmic webs such as $\beta$-skeleton (Fang et al., 2019), FoF (Davis et al., 1985), density-based techniques (Klypin and Holtzman, 1997; Springel et al., 2001; Knollmann and Knebe, 2009), T-web(Hahn et al., 2007a; Forero-Romero et al., 2009a), and V-web (Hoffman et al., 2012a; Forero-Romero et al., 2014a).
In 2019, the $\beta$-skeleton analysis was employed for the first time to characterize the cosmic web, by analyzing both observational data and accompanying simulations. The observational data utilized the BOSS DR12 CMASS (Reid et al., 2015) galaxy sample, and the BigMultiDark N-body simulation conducted under the Planck 2015 Cosmology (Klypin et al., 2016) served as a test sample. By varying the $\beta$ value and examining the many different scenarios such as redshift evolution and redshift space distortion, it is observed that the number of skeleton connections decreases with increasing $\beta$, resulting in a sparser cosmic web. For instance, when $\beta=1$, numerous connections are generated, irrespective of whether the nodes are located in voids or filaments. However, with setting $\beta=3$, the resulting cosmic web closely resembles that of real observations, indicating a nearly perfect match in the number of connections to the actual galaxies. Finally, at $\beta=10$, only small and relatively isolated compact groups of galaxies are identified and connected.
@@
Recently, the mark weighted correlation function (MCF) White (2016)), which assigns density weights to different features of galaxies to extract non-Gaussian information on LSS, has proven effective in capturing more detailed clustering information. MCFs are computationally more feasible compared to other statistical methods capable of extracting non-Gaussian information from LSS. By selecting weights based on the $\alpha$-th power of the local density, the weighting amplifies either dense or low-density regions. This method can enhance the distinction between galaxy clusters in dense and sparse regions, resulting in more stringent constraints compared with 2PCF. Yang et al. (2020a); Lai et al. (2023) demonstrate that the joint constraint of the MCFs with different $\alpha$ values significantly improved constraints on cosmological parameters such as $\Omega_{m}$ and $w$.  Satpathy et al. (2019) presents measurements of MCFs of LOWZ galaxies from SDSS BOSS DR12 to distinguish between the $\Lambda$CDM model and $f(R)$ gravity models. Xiao et al. (2022) employed density gradients as weights, and find that the gradient-weighting scheme would produce more robust parameter constraints compared to the density-marked scheme. Furthermore, combining these two schemes yields even stronger constraints than using either one alone.
In this study, we present the first application of the $\beta$-skeleton as weights in MCFs, serving as a statistical measure of information about the environment contained in the knots of the $\beta$-skeleton. Constructing undirected graph information from the $\beta$-skeleton allows for extracting additional information about LSS. We will apply three weighting schemes to MCFs utilizing undirected graph information derived from the $\beta$-skeleton. These schemes involve weighting by: i) the number of neighboring galaxies around each galaxy, denoted as $N_{\rm con}$; ii) the average distance of each galaxy from its surrounding neighbors, denoted as $\bar{D}_{\rm nei}$; iii) the reciprocal of the average distance of each galaxy from its surrounding neighbors, $1/\bar{D}_{\rm nei}$.
This paper is organized as follows: In Sect. II, we provide a brief introduction to the definition and basic properties of the $\beta$-skeleton, outlining the methodology for calculating the MCFs on the data. Sect. III details the COLA and PATCHY simulations, as well as the Sloan Digital Sky Survey III (SDSS-III) Baryon Oscillation Spectroscopic Survey (BOSS) Data Release 12 (DR12) galaxy sample, which are utilized in the subsequent analysis. In Sect. IV, we present our findings by applying the MCFs using the weights derived from the $\beta$-skeleton quantities. Finally, we summarize our results and draw conclusions in Sect. V.
@@ II Data
To investigate the sensitivity of MCF statistics using the properties of the $\beta$-skeleton as the weighting scheme for cosmological parameters, we need to compare actual observations with a series of simulations. Our analysis relies on observations of SDSS BOSS DR12 CMASS galaxies. Additionally, the analysis in this work depends on a set of fast simulations generated using the COmoving Lagrangian Acceleration (COLA) method (Tassev et al., 2013a). The galaxy mock catalog accurately reproduces the statistical properties of observed CMASS galaxies by employing a recently proposed scheme Ding et al. (2024) that utilizes the subhalo abundance-matching (SHAM) procedure and involves adjusting the COLA simulation parameters. Additionally, we utilize the Multidark PATCHY simulations (referred to as PATCHY) Kitaura et al. (2016a) to estimate the relevant covariance.
@@ II.1 Observational data
The inclusion of the Baryon Oscillation Spectroscopic Survey (BOSS) is a component of SDSS-III (Eisenstein et al., 2011; Bolton et al., 2012; Dawson et al., 2012). The detection of the characteristic scale imprinted by baryon acoustic oscillations (BAO) in the early Universe is aimed for by the measurement of the spatial distribution of luminous red galaxies (LRGs) and quasars through BOSS (Eisenstein et al., 2001).
@@
II.2 COLA Simulations and mock galaxy catalogues
COLA [^1] https://bitbucket.org/tassev/colacode/src/hg/, a hybrid approach integrating second-order LPT (2LPT) and N-body algorithms, proves to be an effective solution for simulating dark matter (DM) particles. Perturbation theory has demonstrated success in describing large scales, allowing the linear growth rate to replace time integration in N-body simulations. COLA leverages this by employing a comoving frame, with observers following trajectories calculated in perturbation theory. Importantly, COLA accurately calculates large-scale dynamics using 2LPT while assigning the resolution of small scales to the N-body code, without demanding on an exact representation of the internal dynamics of halos. Consequently, COLA can trade accuracy at small scales for computational speed without compromising accuracy at large scales.
@@ II.3 PATCHY Simulation for Convariance Estimation
To accurately estimate the covariance matrix for MCFs, 1000 PATCHY catalogs (Kitaura et al., 2016b) are employed in this study. The PATCHY (PerturbAtion Theory Catalog generator of Halo and galaxY distributions) mock employs a method that integrates an efficient structure formation model with a local, non-linear, scale-dependent, and stochastic biasing scheme to produce mock halo catalogues. Augmented Lagrangian Perturbation Theory is utilized for simulating the structure formation, combining 2LPT on large scales with the spherical collapse model on smaller scales. This technique generates a dark matter density field on a mesh starting from Gaussian fluctuations and calculates the peculiar velocity field. In short, PATCHY mock catalogs are generated using approximate gravity solvers and analytical-statistical biasing models, with halo occupation parameters adjusted such that the mocks well reproduce the BOSS two- and three-point statistics. They have been calibrated to the BigMultiDark N-body simulation with high resolution  (Kitaura et al., 2014; Rodríguez-Torres et al., 2016a), which utilizes $3840^{3}$ particles in a volume of $(2.5~{}h^{-1}\,\rm{Gpc})^{3}$. This simulation assumes a $\Lambda$CDM cosmology with $\Omega_{m}=0.307$, $\Omega_{b}=0.048$, $\sigma_{8}=0.82$, $n_{s}=0.96$, and $h=0.67$.
Figure 1: Comparison of the spatial and redshift distributions for observational galaxies and simulation samples. Upper: the R.A.–Dec. distribution of CMASS galaxies in redshift space is illustrated for the range of $z\in[0.45,0.55]$. Middle: the R.A.–Dec. distribution of generated COLA mock halos for $\Omega_{m}=0.31$, showing the same angular distribution as CMASS galaxies. Bottom: the histograms of the number of galaxies (divided into 150 $z$-bins) are presented for CMASS (red solid) and COLA simulations with $\Omega_{m}=0.31$ (blue dashed-dotted), 0.25 (orange dashed), and 0.4 (green dotted), respectively. As seen, the redshift distributions of the three COLA samples closely match that of CMASS.
@@ III.1 $\beta$-skeleton analysis
We will perform the $\beta$-skeleton analysis and use its properties as weights to conduct the MCFs for CMASS data and COLA mocks. Let us first briefly summarize the $\beta$-skeleton theory and explain how it is used to characterize the cosmic web.
Stemming from the fields of computational geometry and geometric graph theory, the $\beta$-skeleton algorithm has been widely applied in image analysis, machine learning, visual perception, and pattern recognition Edelsbrunner et al. (1983); Amenta et al. (1998a); Zhang and King (2002). In the realm of web finders, a class of algorithms, including the $\beta$-skeleton, is employed to construct a graph describing the degree of connectedness, starting from a set of 3D spatial points. This process, resembling the minimum spanning tree (MST) algorithm Barrow et al. (1985), has a key distinction – the resulting graph depends on the continuous $\beta$ parameter. Additionally, web finders designed based on topological persistence, such as DisPerSE Sousbie (2011a), Betti numbers Pranav et al. (2019a) , the caustic skeleton defined based on Lagrangian fluid dynamics Feldbrugge et al. (2017), and others, are related to the $\beta$-skeleton. For more detailed information regarding the $\beta$-skeleton in topology and geometric graph theory, please refer to Kirkpatrick and Radke (1985); Correa and Lindstrom (2013).
@@
The $\beta$-skeleton, as defined above, possesses various interesting mathematical properties. As $\beta$ continuously varies from 0 to $\infty$, the constructed graphs transition from a complete graph to an empty graph. The special case of $\beta=1$ results in the Gabriel graph, known to include the Euclidean MST. In image analysis, for instance, it has been utilized to reconstruct the shape of a two-dimensional object based on a set of sample points on the object’s boundary. It has been proven that the Circle-based graphs with $\beta=1.7$ can accurately reconstruct smooth surface boundaries without generating extra edges when samples are dense enough relative to local curvature. It is apparent that the volume of the excluded region exhibits a monotonically increasing relationship with $\beta$. As the excluded region expands, the requirements become more stringent and challenging, making it less likely to find connected paired galaxies. The $\beta$-skeleton code used in this paper is available at the following link [^2] https://github.com/xiaodongli1986/LSSCode.
The $\beta$-skeleton algorithm is applied with $\beta$ values of 1, 3, and 5, respectively, to both CMASS galaxies and COLA mock samples for various values of $\Omega_{m}$. The derived cosmic webs are depicted in Fig. 2. For comparison, the Friend-of-Friend (FoF) algorithm is applied with linking lengths (denoted as $L_{\rm link}$) set to 10 and 15 $h^{-1}{\rm Mpc}$, and the resulting FoF cosmic webs are also displayed (bottom). All maps are selected within the range of R.A. $\in[200^{\circ},208^{\circ}]$ and Dec. $\in[30^{\circ},38^{\circ}]$, corresponding to a slice of side length $183.96\times 183.96$ $h^{-1}{\rm Mpc}$, respectively. The depth of these maps is 227.54 $h^{-1}{\rm Mpc}$, corresponding to the redshift range of $z\in[0.45,0.55]$.
@@
Moreover, filament-like structures are generated from the galaxy sample by the $\beta$-skeleton; this is most clearly detected when $\beta=3$. More densely and sparsely distributed filaments are produced by larger and smaller values of $\beta$, unlike cosmic webs observed in reality. Additionally, “knots” within the structure are formed by some of these galaxies, where three or more galaxies are connected. Moreover, there are isolated structures with a relatively small number of group members. Varying $\beta$ strongly influences the overall shape of the skeleton maps. For instance, when $\beta=1$, it roughly corresponds to computing two-point correlations, resulting in numerous connections being generated, irrespective of whether these connections lie within a filament or not. With $\beta=3$, the generated set of structures closely approximates the observed cosmic web, with the number of connections comparable to the actual number of galaxies. In contrast, when $\beta=5$, connections are sparse, as expected, identifying and connecting only the small and relatively isolated compact groups of galaxies. Moreover, statistically, the connection length, denoted as $L_{\rm con}$, varies with $\beta$. As $\beta$ increases, $L_{\rm con}$ tends to decrease and appears more concentrated. This is because the stringent threshold makes connecting two points separated by a large distance more challenging. Also, upon visual inspection, it may be challenging to identify obvious differences in the skeleton graphs corresponding to different values of $\Omega_{m}$.
@@
Figure 2: Visualization of cosmic webs produced by the $\beta$-skeleton and Friend-of-Friend (FoF) algorithms for comparison. All maps are selected within the range of R.A. $\in[200^{\circ},208^{\circ}]$ and Dec. $\in[30^{\circ},38^{\circ}]$. In the visualization, the scatter represents galaxies, and the resulting connections between galaxy pairs are also depicted. The color is indicative of the co-moving distance of the respective galaxies. Upper: the $\beta$-skeleton web of CMASS-NGC galaxies is illustrated for $\beta=1$ (left) and $\beta=5$ (right). The 2nd and 3rd rows: the $\beta$-skeleton web is obtained by choosing $\beta=3$ for CMASS-NGC and COLA simulations with $\Omega_{m}=0.31$, $\Omega_{m}=0.25$, $\Omega_{m}=0.4$. Bottom: the cosmic web obtained from FoF with link lengths, $L_{\rm link}$, set to $10$ (left) and $15$ (right), respectively. In comparison to $L_{\rm link}=$ 10 and 15 in FoF, the $\beta$-skeleton is adaptive to both low and high-density regions. As observed, the $\beta$-skeleton graph is more reliable and natural compared to the cosmic webs produced by FoF, highlighting the advantages of the $\beta$-skeleton graph. $\beta=3$ is found to be optimal and better corresponds to the real cosmic web, represented as the plot of $\Omega_{m}=0.31,\beta=3$, in contrast to the two graphs that exhibit overly dense ($\beta=1$) or overly sparse ($\beta=5$) patterns.

### v3

@@ Abstract
The $\beta$-skeleton approach can be conveniently utilized to construct the cosmic web based on the spatial geometry distribution of galaxies, particularly in sparse samples. This method plays a key role in establishing the three-dimensional structure of the Universe and serves as a tool for quantitatively characterizing the nature of the cosmic web. This study is the first application of $\beta$-skeleton information as weights in mark weighted correlation functions (MCFs), presenting a novel statistical measure. We have applied the $\beta$-skeleton approach to the CMASS NGC galaxy samples from SDSS BOSS DR12 in the redshift interval $0.45\leq z\leq 0.55$. Additionally, we applied this approach to three COLA cosmological simulations with different settings ($\Omega_{m}=0.25,\Omega_{m}=0.31,\Omega_{m}=0.4$) for comparison. We measured three MCFs, each weighted by i) the number of neighboring galaxies around each galaxy, ii) the average distance of each galaxy from its surrounding neighbors, and iii) the reciprocal of the average distance of each galaxy from its surrounding neighbors. By comparing measurements and calculating corresponding $\chi^{2}$ statistics, we observe high sensitivity to the cosmological parameter $\Omega_{m}$ through a joint analysis of the two-point correlation and three MCFs.
@@ I Introduction
According to the principles of cosmology, the Universe is considered homogeneous and isotropic at a large scale. However, the combined influence of gravity and the ongoing expansion of the Universe results in a complex and nonlinear structure. In terms of spatial configuration, the unique and intricate web structure formed by the anisotropic collapse of gravity is referred to as the “cosmic web” (Bardeen et al., 1986; de Lapparent et al., 1986; Huchra et al., 2012; Tegmark et al., 2004; Guzzo et al., 2014; Suárez-Pérez et al., 2021). Topologically, based on the dimensions of gravitational collapse, the cosmic web is typically categorized into knots, filaments, sheets, and voids.
The study of the structure of the cosmic web relies on large-scale redshift surveys, and currently, there are a number of large-scale observational projects, such as the Sloan Digital Sky Survey (SDSS) (York et al., 2000; Eisenstein et al., 2005; Percival et al., 2007; Anderson et al., 2012; Sånchez et al., 2012, 2013; Anderson et al., 2014; Ross et al., 2015; Beutler et al., 2017; Sånchez et al., 2017; Alam et al., 2017; Chuang et al., 2017), the 2-degree Field Galaxy Redshift Survey (2dFGRS) (Colless et al., 2001), the 6-degree Field Galaxy Redshift Survey (6dFGRS) (Beutler et al., 2012, 2011), and the WiggleZ Dark Energy Survey (Parkinson et al., 2012).
The ongoing stage-IV surveys, such as the Dark Energy Spectroscopic Instrument (DESI) DESI Collaboration et al. (2016, 2022), are expected to further deepen our understanding of the Universe. They possess the capability to observe wider and deeper fields of view with higher resolution, allowing for the collection of more accurate and extensive data and providing additional information about the nonlinear scale structure of the Universe.
@@
The exploration of the cosmic web has consistently been a significant topic in the study of the Universe. Since early maps of the Universe confirmed the existence of such webs through galaxy redshift measurements, there has been a continuous effort over the past 40 years to find a consistent and stable method to define these weblike structures. Currently, the commonly used analytical tools for studying cosmic webs such as $\beta$-skeleton (Fang et al., 2019), FoF (Davis et al., 1985), density-based techniques (Klypin and Holtzman, 1997; Springel et al., 2001a; Knollmann and Knebe, 2009), T-web (Hahn et al., 2007; Forero-Romero et al., 2009), and V-web (Hoffman et al., 2012; Forero-Romero et al., 2014).
In 2019, the $\beta$-skeleton analysis was employed for the first time to characterize the cosmic web, by analyzing both observational data and accompanying simulations. The observational data utilized the BOSS DR12 CMASS (Reid et al., 2016) galaxy sample, and the BigMultiDark N-body simulation conducted under the Planck 2015 Cosmology (Klypin et al., 2016) served as a test sample. By varying the $\beta$ value and examining the many different scenarios such as redshift evolution and redshift space distortion, it is observed that the number of skeleton connections decreases with increasing $\beta$, resulting in a sparser cosmic web. For instance, when $\beta=1$, numerous connections are generated, irrespective of whether the nodes are located in voids or filaments. However, with setting $\beta=3$, the resulting cosmic web closely resembles that of real observations, indicating a nearly perfect match in the number of connections to the actual galaxies. Finally, at $\beta=10$, only small and relatively isolated compact groups of galaxies are identified and connected.
@@
Recently, the mark weighted correlation function (MCF) White (2016)), which assigns density weights to different features of galaxies to extract non-Gaussian information on LSS, has proven effective in capturing more detailed clustering information. MCFs are computationally more feasible compared to other statistical methods capable of extracting non-Gaussian information from LSS. By selecting weights based on the $\alpha$th power of the local density, the weighting amplifies either dense or low-density regions. This method can enhance the distinction between galaxy clusters in dense and sparse regions, resulting in more stringent constraints compared with 2PCF. References Yang et al. (2020); Lai et al. (2023) demonstrate that the joint constraint of the MCFs with different $\alpha$ values significantly improved constraints on cosmological parameters such as $\Omega_{m}$ and $w$. Reference Satpathy et al. (2019) presents measurements of MCFs of LOWZ galaxies from SDSS BOSS DR12 to distinguish between the $\Lambda$CDM model and $f(R)$ gravity models. Reference Xiao et al. (2022) employed density gradients as weights, and find that the gradient-weighting scheme would produce more robust parameter constraints compared to the density-marked scheme. Furthermore, combining these two schemes yields even stronger constraints than using either one alone.
In this study, we present the first application of the $\beta$-skeleton as weights in MCFs, serving as a statistical measure of information about the environment contained in the knots of the $\beta$-skeleton. Constructing undirected graph information from the $\beta$-skeleton allows for extracting additional information about LSS. We will apply three weighting schemes to MCFs utilizing undirected graph information derived from the $\beta$-skeleton. These schemes involve weighting by i) the number of neighboring galaxies around each galaxy, denoted as $N_{\rm con}$; ii) the average distance of each galaxy from its surrounding neighbors, denoted as $\bar{D}_{\rm nei}$; and iii) the reciprocal of the average distance of each galaxy from its surrounding neighbors, $1/\bar{D}_{\rm nei}$.
This paper is organized as follows: Sect. II details the COLA and PATCHY simulations, as well as the Sloan Digital Sky Survey III (SDSS-III) Baryon Oscillation Spectroscopic Survey (BOSS) Data Release 12 (DR12) galaxy sample, which are utilized in the subsequent analysis. In Sec. III, we provide a brief introduction to the definition and basic properties of the $\beta$-skeleton, outlining the methodology for calculating the MCFs on the data. In Sec. IV, we present our findings by applying the MCFs using the weights derived from the $\beta$-skeleton quantities. Finally, we summarize our results and draw conclusions in Sec. V.
@@ II Data
To investigate the sensitivity of MCF statistics using the properties of the $\beta$-skeleton as the weighting scheme for cosmological parameters, we need to compare actual observations with a series of simulations. Our analysis relies on observations of SDSS BOSS DR12 CMASS galaxies. Additionally, the analysis in this work depends on a set of fast simulations generated using the COmoving Lagrangian Acceleration (COLA) method (Tassev et al., 2013). The galaxy mock catalog accurately reproduces the statistical properties of observed CMASS galaxies by employing a recently proposed scheme Ding et al. (2024) that utilizes the subhalo abundance-matching (SHAM) procedure and involves adjusting the COLA simulation parameters. Note that to ensure the accuracy of COLA for our analysis, we have compared various statistics with those derived from GADGET simulations using the same cosmological parameters, and found agreement between the two (see the Appendix for details). Additionally, we utilize the Multidark PATCHY simulations (referred to as PATCHY) Kitaura et al. (2016a) to estimate the relevant covariance.
@@ II.1 Observational data
The inclusion of the Baryon Oscillation Spectroscopic Survey (BOSS) is a component of SDSS-III (Eisenstein et al., 2011; Bolton et al., 2012; Dawson et al., 2013). The detection of the characteristic scale imprinted by baryon acoustic oscillations (BAO) in the early Universe is aimed for by the measurement of the spatial distribution of luminous red galaxies (LRGs) and quasars through BOSS (Eisenstein et al., 2001).
@@
II.2 COLA simulations and mock galaxy catalogs
COLA, [^1] https://bitbucket.org/tassev/colacode/src/hg/ a hybrid approach integrating second-order LPT (2LPT) and N-body algorithms, proves to be an effective solution for simulating dark matter (DM) particles. Perturbation theory has demonstrated success in describing large scales, allowing the linear growth rate to replace time integration in N-body simulations. COLA leverages this by employing a comoving frame, with observers following trajectories calculated in perturbation theory. Importantly, COLA accurately calculates large-scale dynamics using 2LPT while assigning the resolution of small scales to the N-body code, without demanding on an exact representation of the internal dynamics of halos. Consequently, COLA can trade accuracy at small scales for computational speed without compromising accuracy at large scales.
@@ II.3 PATCHY simulation for convariance estimation
To accurately estimate the covariance matrix for MCFs, 1000 PATCHY catalogs (Kitaura et al., 2016b) are employed in this study. The PATCHY (PerturbAtion Theory Catalog generator of Halo and galaxY distributions) mock employs a method that integrates an efficient structure formation model with a local, nonlinear, scale-dependent, and stochastic biasing scheme to produce mock halo catalogs. Augmented Lagrangian Perturbation Theory is utilized for simulating the structure formation, combining 2LPT on large scales with the spherical collapse model on smaller scales. This technique generates a dark matter density field on a mesh starting from Gaussian fluctuations and calculates the peculiar velocity field. In short, PATCHY mock catalogs are generated using approximate gravity solvers and analytical-statistical biasing models, with halo occupation parameters adjusted such that the mocks well reproduce the BOSS two- and three-point statistics. They have been calibrated to the BigMultiDark N-body simulation with high resolution  (Kitaura et al., 2014; Rodríguez-Torres et al., 2016), which utilizes $3840^{3}$ particles in a volume of $(2.5~{}h^{-1}\,\rm{Gpc})^{3}$. This simulation assumes a $\Lambda$CDM cosmology with $\Omega_{m}=0.307$, $\Omega_{b}=0.048$, $\sigma_{8}=0.82$, $n_{s}=0.96$, and $h=0.67$.
Figure 1: Comparison of the spatial and redshift distributions for observational galaxies and simulation samples. Upper: the R.A.–Dec. distribution of CMASS-NGC galaxies in redshift space is illustrated for the range of $z\in[0.45,0.55]$. Middle: the R.A.–Dec. distribution of generated COLA mock halos for $\Omega_{m}=0.31$, showing the same angular distribution as CMASS-NGC galaxies. Bottom: the histograms of the number of galaxies (divided into 150 $z$-bins) are presented for CMASS-NGC (red solid) and COLA simulations with $\Omega_{m}=0.31$ (blue dashed-dotted), 0.25 (orange dashed), and 0.4 (green dotted), respectively. As seen, the redshift distributions of the three COLA samples closely match that of CMASS-NGC.
@@ III.1 $\beta$-skeleton analysis
We will perform the $\beta$-skeleton analysis and use its properties as weights to conduct the MCFs for CMASS-NGC data and COLA mocks. Let us first briefly summarize the $\beta$-skeleton theory and explain how it is used to characterize the cosmic web.
Stemming from the fields of computational geometry and geometric graph theory, the $\beta$-skeleton algorithm has been widely applied in image analysis, machine learning, visual perception, and pattern recognition Edelsbrunner et al. (1983); Amenta et al. (1998); Zhang and King (2002). In the realm of web finders, a class of algorithms, including the $\beta$-skeleton, is employed to construct a graph describing the degree of connectedness, starting from a set of 3D spatial points. This process, resembling the minimum spanning tree (MST) algorithm Barrow et al. (1985), has a key distinction – the resulting graph depends on the continuous $\beta$ parameter. Additionally, web finders designed based on topological persistence, such as DisPerSE Sousbie (2011), Betti numbers Pranav et al. (2019) , the caustic skeleton defined based on Lagrangian fluid dynamics Feldbrugge et al. (2018), and others, are related to the $\beta$-skeleton. For more detailed information regarding the $\beta$-skeleton in topology and geometric graph theory, please refer to Kirkpatrick and Radke (1985); Correa and Lindstrom (2013).
@@
The $\beta$-skeleton, as defined above, possesses various interesting mathematical properties. As $\beta$ continuously varies from 0 to $\infty$, the constructed graphs transition from a complete graph to an empty graph. The special case of $\beta=1$ results in the Gabriel graph, known to include the Euclidean MST. In image analysis, for instance, it has been utilized to reconstruct the shape of a two-dimensional object based on a set of sample points on the object’s boundary. It has been proven that the Circle-based graphs with $\beta=1.7$ can accurately reconstruct smooth surface boundaries without generating extra edges when samples are dense enough relative to local curvature. It is apparent that the volume of the excluded region exhibits a monotonically increasing relationship with $\beta$. As the excluded region expands, the requirements become more stringent and challenging, making it less likely to find connected paired galaxies. The $\beta$-skeleton code used in this paper is available at the following link. [^2] https://github.com/xiaodongli1986/LSSCode
The $\beta$-skeleton algorithm is applied with $\beta$ values of 1, 3, and 5, respectively, to both CMASS-NGC galaxies and COLA mock samples for various values of $\Omega_{m}$. The derived cosmic webs are depicted in Fig. 2. For comparison, the Friend-of-Friend (FoF) algorithm is applied with linking lengths (denoted as $L_{\rm link}$) set to 10 and 15 $h^{-1}{\rm Mpc}$, and the resulting FoF cosmic webs are also displayed (bottom). All maps are selected within the range of R.A. ∈$\in[200^{\circ},208^{\circ}]$ and Dec. ∈$\in[30^{\circ},38^{\circ}]$, corresponding to a slice of side length $183.96\times 183.96$ $h^{-1}{\rm Mpc}$, respectively. The depth of these maps is 227.54 $h^{-1}{\rm Mpc}$, corresponding to the redshift range of $z\in[0.45,0.55]$.
@@
Moreover, filamentlike structures are generated from the galaxy sample by the $\beta$-skeleton; this is most clearly detected when $\beta=3$. More densely and sparsely distributed filaments are produced by larger and smaller values of $\beta$, unlike cosmic webs observed in reality. Additionally, “knots” within the structure are formed by some of these galaxies, where three or more galaxies are connected. Moreover, there are isolated structures with a relatively small number of group members. Varying $\beta$ strongly influences the overall shape of the skeleton maps. For instance, when $\beta=1$, it roughly corresponds to computing two-point correlations, resulting in numerous connections being generated, irrespective of whether these connections lie within a filament or not. With $\beta=3$, the generated set of structures closely approximates the observed cosmic web, with the number of connections comparable to the actual number of galaxies. In contrast, when $\beta=5$, connections are sparse, as expected, identifying and connecting only the small and relatively isolated compact groups of galaxies. Moreover, statistically, the connection length, denoted as $L_{\rm con}$, varies with $\beta$. As $\beta$ increases, $L_{\rm con}$ tends to decrease and appears more concentrated. This is because the stringent threshold makes connecting two points separated by a large distance more challenging. Also, upon visual inspection, it may be challenging to identify obvious differences in the skeleton graphs corresponding to different values of $\Omega_{m}$.
@@
Figure 2: Visualization of cosmic webs produced by the $\beta$-skeleton and Friend-of-Friend (FoF) algorithms for comparison. All maps are selected within the range of R.A. ∈$\in[200^{\circ},208^{\circ}]$ and Dec. ∈$\in[30^{\circ},38^{\circ}]$. In the visualization, the scatter represents galaxies, and the resulting connections between galaxy pairs are also depicted. The color is indicative of the comoving distance of the respective galaxies. Upper: the $\beta$-skeleton web of CMASS-NGC galaxies is illustrated for $\beta=1$ (left) and $\beta=5$ (right). The second and third rows: the $\beta$-skeleton web is obtained by choosing $\beta=3$ for CMASS-NGC and COLA simulations with $\Omega_{m}=0.31$, $\Omega_{m}=0.25$, $\Omega_{m}=0.4$. Bottom: the cosmic web obtained from FoF with link lengths, $L_{\rm link}$, set to $10$ (left) and $15$ (right), respectively. In comparison to $L_{\rm link}=$ 10 and 15 in FoF, the $\beta$-skeleton is adaptive to both low and high-density regions. As observed, the $\beta$-skeleton graph is more reliable and natural compared to the cosmic webs produced by FoF, highlighting the advantages of the $\beta$-skeleton graph. $\beta=3$ is found to be optimal and better corresponds to the real cosmic web, represented as the plot of $\Omega_{m}=0.31,\beta=3$, in contrast to the two graphs that exhibit overly dense ($\beta=1$) or overly sparse ($\beta=5$) patterns.